OpenStudy (anonymous):

\sqrt[3]{125}-2\sqrt[4]{81\div16}

OpenStudy (anonymous):

\[\sqrt[3]{125}-2\sqrt[4]{\frac{81}{16}}\] \[5-\cancel2(\frac{3}{\cancel2})\] \[5-3=2\]

OpenStudy (anonymous):

answer is 2